A body intially at rest breaks into two pieces having mass 2M and 3M respectively together having total kinetic emergy E . the piece of mass 2M after having kinetic energy

 A  body intially at rest breaks into two pieces having mass 2M and 3M  respectively together having total kinetic emergy E . the piece of mass 2M after having kinetic energy
 

Grade:12th pass

1 Answers

Arun
6 years ago
 
let speed of 2m,3m piece be u,v respectively
by conservation of momentum
2mu=3mv
u=3v/2
E=1/2 (2m)u^2 + 1/2 (3m)v^2
put u=3v/2
E=9/4mv^2+3/2mv^2
E=15/4mv^2
ke of 2kg block = 1 /2 (2m)u^2=9/4mv^2
=9(E/15)
=3E/5